OpenStudy (anonymous):

[7.07] Choose the correct product of (8x − 4)(8x + 4

OpenStudy (dan815):

go to youtube look up foil method

OpenStudy (anonymous):

dan is correct....this requires the FOIL method. Let me explain.... F -- first numbers O -- outside numbers I -- inside numbers L -- last numbers (8x - 4)(8x + 4) F -- multiply the first numbers in each set....8x * 8x = 64x^2 O -- multiply outer numbers in each set ...8x * 4 = 32x I -- multiply inside numbers of each set ....-4 * 8x = -32x L -- multiply last numbers in each set ... -4 * 4 = - 16 Now lets put down what we got... 64x^2 + 32x - 32x - 16 (combine like terms) 64x^2 - 16 = 8(8x^2 - 2)